23 × ( – 99) = ___ × ( – 100 + ___ ) = 23 × ___ + 23 × ___

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

23, 1, -100, 1:
23 × (–99) = 23 × (–100 + 1 )
= 23 × (-100) + 23 × 1 (Distributive property of integers)

Thanks